# Efficient Model-Stealing Attacks Against Inductive Graph Neural Networks

### v2

@@ Appendix D Additional result - Type II attack
Accuracy Fidelity Accuracy Fidelity Accuracy Fidelity
Shen et al. 0.736$\pm$0.000 0.855$\pm$0.000 0.733$\pm$0.001 0.845$\pm$0.000 0.747$\pm$0.000 0.856$\pm$0.000
Prediction ours 0.745$\pm$0.004 0.811$\pm$0.001 0.742$\pm$0.001 0.820$\pm$0.003 0.749$\pm$0.002 0.815$\pm$0.002
Shen et al. 0.668$\pm$0.000 0.746$\pm$0.000 0.704$\pm$0.002 0.794$\pm$0.002 0.636$\pm$0.002 0.697$\pm$0.003
Projection ours 0.701$\pm$0.003 0.756$\pm$0.004 0.715$\pm$0.003 0.775$\pm$0.004 0.688$\pm$0.004 0.742$\pm$0.006
Shen et al. 0.733$\pm$0.001 0.837$\pm$0.000 0.746$\pm$0.000 0.858$\pm$0.000 0.748$\pm$0.005 0.861$\pm$0.010
DBLP Embedding ours 0.736$\pm$0.004 0.790$\pm$0.003 0.745$\pm$0.008 0.778$\pm$0.015 0.705$\pm$0.001 0.742$\pm$0.002
Shen et al. 0.835$\pm$0.000 0.911$\pm$0.000 0.817$\pm$0.001 0.910$\pm$0.007 0.832$\pm$0.000 0.913$\pm$0.000 0
Prediction ours 0.850$\pm$0.000 0.875$\pm$0.001 0.826$\pm$0.007 0.880$\pm$0.007 0.849$\pm$0.001 0.875$\pm$0.001
Shen et al. 0.749$\pm$0.000 0.800$\pm$0.000 0.797$\pm$0.001 0.892$\pm$0.001 0.751$\pm$0.000 0.806$\pm$0.000
Projection ours 0.838$\pm$0.002 0.867$\pm$0.001 0.826$\pm$0.005 0.881$\pm$0.005 0.837$\pm$0.000 0.860$\pm$0.001
Shen et al. 0.844$\pm$0.000 0.911$\pm$0.000 0.827$\pm$0.000 0.911$\pm$0.000 0.838$\pm$0.001 0.931$\pm$0.00
Pubmed Embedding ours 0.851$\pm$0.002 0.867$\pm$0.002 0.841$\pm$0.000 0.874$\pm$0.001 0.845$\pm$0.002 0.847$\pm$0.002
Shen et al. 0.822$\pm$0.001 0.902$\pm$0.000 0.835$\pm$0.001 0.893$\pm$0.001 0.828$\pm$0.000 0.895$\pm$0.001
Prediction ours 0.817$\pm$0.004 0.849$\pm$0.002 0.843$\pm$0.003 0.880$\pm$0.002 0.824$\pm$0.004 0.862$\pm$0.001
Shen et al. 0.682$\pm$0.008 0.713$\pm$0.000 0.742$\pm$0.003 0.787$\pm$0.002 0.624$\pm$0.021 0.661$\pm$0.023
Projection ours 0.730$\pm$0.001 0.749$\pm$0.001 0.757$\pm$0.001 0.783$\pm$0.001 0.695$\pm$0.007 0.712$\pm$0.001
Shen et al. 0.825$\pm$0.001 0.909$\pm$0.000 0.833$\pm$0.001 0.918$\pm$0.002 0.819$\pm$0.001 0.892$\pm$0.001
Citeseer Embedding ours 0.814$\pm$0.006 0.843$\pm$0.005 0.844$\pm$0.001 0.869$\pm$0.006 0.812$\pm$0.005 0.820$\pm$0.005
Shen et al. 0.907$\pm$0.000 0.922$\pm$0.000 0.905$\pm$0.002 0.930$\pm$0.004 0.886$\pm$0.003 0.906$\pm$0.004
Prediction ours 0.922$\pm$0.001 0.932$\pm$0.004 0.896$\pm$0.003 0.925$\pm$0.004 0.879$\pm$0.011 0.895$\pm$0.011
Shen et al. 0.446$\pm$0.002 0.444$\pm$0.000 0.395$\pm$0.091 0.396$\pm$0.086 0.500$\pm$0.002 0.495$\pm$0.002
Projection ours 0.500$\pm$0.006 0.490$\pm$0.006 0.497$\pm$0.008 0.492$\pm$0.008 0.400$\pm$0.075 0.399$\pm$0.071
Shen et al. 0.750$\pm$0.011 0.766$\pm$0.000 0.864$\pm$0.026 0.878$\pm$0.032 0.589$\pm$0.026 0.606$\pm$0.025
Amazon Embedding ours 0.906$\pm$0.005 0.924$\pm$0.003 0.861$\pm$0.032 0.882$\pm$0.034 0.839$\pm$0.011 0.851$\pm$0.012
Shen et al. 0.948$\pm$ 0.001 0.960$\pm$ 0.002 0.906$\pm$ 0.001 0.927$\pm$ 0.002 0.939$\pm$ 0.003 0.953$\pm$ 0.002
Prediction ours 0.948$\pm$0.002 0.959$\pm$0.003 0.895$\pm$0.006 0.911$\pm$0.007 0.944$\pm$0.001 0.956$\pm$0.001
Shen et al. 0.821$\pm$ 0.001 0.827$\pm$ 0.001 0.859$\pm$ 0.002 0.937$\pm$ 0.002 0.799$\pm$ 0.021 0.850$\pm$ 0.031
Projection ours 0.789$\pm$0.006 0.795$\pm$0.007 0.673$\pm$0.020 0.684$\pm$0.021 0.758$\pm$0.009 0.765$\pm$0.008
Shen et al. 0.945$\pm$ 0.001 0.826$\pm$ 0.001 0.920$\pm$ 0.002 0.937$\pm$ 0.001 0.930$\pm$ 0.007 0.947$\pm$ 0.011
Coauthor Embedding ours 0.941$\pm$0.002 0.952$\pm$0.001 0.904$\pm$0.009 0.921$\pm$0.008 0.939$\pm$0.001 0.948$\pm$0.001
Shen et al. 0.898$\pm$0.000 0.947$\pm$0.000 0.900$\pm$0.000 0.958$\pm$0.000 0.892$\pm$0.000 0.936$\pm$0.000
Prediction ours 0.903$\pm$0.004 0.923$\pm$0.005 0.896$\pm$0.007 0.940$\pm$0.003 0.889$\pm$0.012 0.923$\pm$0.007
Shen et al. 0.876$\pm$0.000 0.915$\pm$0.000 0.857$\pm$0.000 0.913$\pm$0.000 0.888$\pm$0.000 0.929$\pm$0.000
Projection ours 0.880$\pm$0.002 0.906$\pm$0.002 0.842$\pm$0.012 0.891$\pm$0.010 0.867$\pm$0.003 0.886$\pm$0.009
Shen et al. 0.900$\pm$0.000 0.927$\pm$0.000 0.845$\pm$0.000 0.880$\pm$0.000 0.839$\pm$0.000 0.855$\pm$0.000
ACM Embedding ours 0.835$\pm$0.010 0.848$\pm$0.016 0.864$\pm$0.013 0.900$\pm$0.011 0.853$\pm$0.003 0.862$\pm$0.007
@@
Accuracy Fidelity Accuracy Fidelity Accuracy Fidelity
Shen et al. 0.761$\pm$0.000 0.904$\pm$0.000 0.759$\pm$0.000 0.864$\pm$0.000 0.764$\pm$0.000 0.891$\pm$0.001
Prediction ours 0.761$\pm$0.002 0.858$\pm$0.002 0.770$\pm$0.007 0.853$\pm$0.005 0.760$\pm$0.005 0.849$\pm$0.008
Shen et al. 0.714$\pm$0.001 0.807$\pm$0.000 0.734$\pm$0.000 0.839$\pm$0.000 0.729$\pm$0.001 0.830$\pm$0.000
Projection ours 0.731$\pm$0.002 0.821$\pm$0.003 0.737$\pm$0.004 0.824$\pm$0.006 0.743$\pm$0.000 0.829$\pm$0.001
Shen et al. 0.757$\pm$0.000 0.887$\pm$0.000 0.776$\pm$0.001 0.885$\pm$0.001 0.762$\pm$0.000 0.895$\pm$0.000
DBLP Embedding ours 0.737$\pm$0.017 0.809$\pm$0.021 0.764$\pm$0.003 0.837$\pm$0.003 0.724$\pm$0.003 0.787$\pm$0.003
Shen et al. 0.855$\pm$0.000 0.956$\pm$0.000 0.839$\pm$0.000 0.944$\pm$0.001 0.856$\pm$0.000 0.952$\pm$0.001
Prediction ours 0.856$\pm$0.002 0.925$\pm$0.001 0.846$\pm$0.003 0.925$\pm$0.003 0.855$\pm$0.001 0.924$\pm$0.000
Shen et al. 0.852$\pm$0.000 0.951$\pm$0.000 0.822$\pm$0.001 0.922$\pm$0.001 0.845$\pm$0.000 0.938$\pm$0.000
Projection ours 0.848$\pm$0.001 0.920$\pm$0.001 0.838$\pm$0.001 0.920$\pm$0.002 0.842$\pm$0.001 0.915$\pm$0.002
Shen et al. 0.859$\pm$0.001 0.958$\pm$0.000 0.842$\pm$0.002 0.944$\pm$0.003 0.848$\pm$0.000 0.953$\pm$0.000
Pubmed Embedding ours 0.859$\pm$0.000 0.920$\pm$0.005 0.848$\pm$0.002 0.924$\pm$0.002 0.847$\pm$0.003 0.895$\pm$0.006
Shen et al. 0.829$\pm$0.001 0.933$\pm$0.000 0.846$\pm$0.001 0.908$\pm$0.000 0.831$\pm$0.000 0.908$\pm$0.000
Prediction ours 0.814$\pm$0.003 0.867$\pm$0.001 0.843$\pm$0.004 0.897$\pm$0.003 0.818$\pm$0.001 0.876$\pm$0.006
Shen et al. 0.685$\pm$0.004 0.727$\pm$0.000 0.736$\pm$0.001 0.792$\pm$0.003 0.719$\pm$0.003 0.756$\pm$0.004
Projection ours 0.718$\pm$0.004 0.762$\pm$0.005 0.750$\pm$0.003 0.797$\pm$0.001 0.666$\pm$0.015 0.704$\pm$0.014
Shen et al. 0.830$\pm$0.001 0.913$\pm$0.000 0.840$\pm$0.001 0.905$\pm$0.001 0.820$\pm$0.002 0.886$\pm$0.001
Citeseer Embedding ours 0.812$\pm$0.005 0.849$\pm$0.004 0.845$\pm$0.002 0.878$\pm$0.001 0.803$\pm$0.004 0.824$\pm$0.002
Shen et al. 0.924$\pm$0.001 0.953$\pm$0.000 0.899$\pm$0.009 0.932$\pm$0.007 0.876$\pm$0.002 0.921$\pm$0.002
Prediction ours 0.922$\pm$0.005 0.932$\pm$0.009 0.909$\pm$0.004 0.930$\pm$0.008 0.899$\pm$0.008 0.913$\pm$0.011
Shen et al. 0.720$\pm$0.004 0.743$\pm$0.000 0.767$\pm$0.011 0.818$\pm$0.008 0.638$\pm$0.002 0.674$\pm$0.002
Projection ours 0.856$\pm$0.013 0.854$\pm$0.014 0.779$\pm$0.010 0.794$\pm$0.004 0.654$\pm$0.028 0.670$\pm$0.033
Shen et al. 0.926$\pm$0.001 0.948$\pm$0.000 0.900$\pm$0.007 0.939$\pm$0.005 0.859$\pm$0.001 0.870$\pm$0.001
Amazon Embedding ours 0.921$\pm$0.003 0.929$\pm$0.003 0.894$\pm$0.018 0.916$\pm$0.035 0.906$\pm$0.004 0.913$\pm$0.005
Shen et al. 0.948$\pm$0.000 0.979$\pm$0.000 0.866$\pm$0.004 0.891$\pm$0.004 0.945$\pm$0.000 0.981$\pm$0.000
Prediction ours 0.946$\pm$0.001 0.967$\pm$0.001 0.884$\pm$0.040 0.909$\pm$0.042 0.947$\pm$0.002 0.971$\pm$0.002
Shen et al. 0.731$\pm$0.001 0.743$\pm$0.000 0.844$\pm$0.007 0.865$\pm$0.008 0.717$\pm$0.005 0.729$\pm$0.005
Projection ours 0.907$\pm$0.011 0.924$\pm$0.011 0.674$\pm$0.053 0.682$\pm$0.054 0.782$\pm$0.003 0.802$\pm$0.005
Shen et al. 0.947$\pm$0.000 0.977$\pm$0.000 0.878$\pm$0.002 0.894$\pm$0.006 0.890$\pm$0.030 0.905$\pm$0.034
Coauthor Embedding ours 0.938$\pm$0.003 0.952$\pm$0.006 0.921$\pm$0.004 0.940$\pm$0.004 0.938$\pm$0.001 0.954$\pm$0.003
Shen et al. 0.907$\pm$0.000 0.950$\pm$0.000 0.902$\pm$0.000 0.950$\pm$0.000 0.905$\pm$0.000 0.943$\pm$0.000
Prediction ours 0.905$\pm$0.003 0.924$\pm$0.002 0.900$\pm$0.006 0.938$\pm$0.002 0.889$\pm$0.006 0.910$\pm$0.012
Shen et al. 0.901$\pm$0.000 0.938$\pm$0.000 0.896$\pm$0.000 0.944$\pm$0.000 0.904$\pm$0.000 0.939$\pm$0.000
Projection ours 0.895$\pm$0.002 0.915$\pm$0.002 0.887$\pm$0.003 0.920$\pm$0.004 0.869$\pm$0.004 0.886$\pm$0.003
Shen et al. 0.907$\pm$0.000 0.952$\pm$0.000 0.868$\pm$0.000 0.919$\pm$0.000 0.870$\pm$0.000 0.895$\pm$0.000
ACM Embedding ours 0.879$\pm$0.004 0.890$\pm$0.007 0.841$\pm$0.027 0.836$\pm$0.033 0.877$\pm$0.007 0.893$\pm$0.006

### v4

@@ Appendix D Additional result - Type II attack
Accuracy Fidelity F1 score Accuracy Fidelity F1 score Accuracy Fidelity F1 score
Shen et al. 0.736$\pm$0.000 0.855$\pm$0.000 0.682$\pm$0.002 0.733$\pm$0.001 0.845$\pm$0.000 0.617$\pm$0.002 0.747$\pm$0.000 0.856$\pm$0.000 0.610$\pm$0.002
Prediction ours 0.745$\pm$0.004 0.811$\pm$0.001 0.631$\pm$0.004 0.742$\pm$0.001 0.820$\pm$0.003 0.637$\pm$0.002 0.749$\pm$0.002 0.815$\pm$0.002 0.620$\pm$0.009
Shen et al. 0.668$\pm$0.000 0.746$\pm$0.000 0.483$\pm$0.002 0.704$\pm$0.002 0.794$\pm$0.002 0.545$\pm$0.002 0.636$\pm$0.002 0.697$\pm$0.003 0.403$\pm$0.001
Projection ours 0.701$\pm$0.003 0.756$\pm$0.004 0.550$\pm$0.005 0.715$\pm$0.003 0.775$\pm$0.004 0.571$\pm$0.005 0.688$\pm$0.004 0.742$\pm$0.006 0.501$\pm$0.003
Shen et al. 0.733$\pm$0.001 0.837$\pm$0.000 0.637$\pm$0.002 0.746$\pm$0.000 0.858$\pm$0.000 0.677$\pm$0.001 0.748$\pm$0.005 0.861$\pm$0.010 0.626$\pm$0.003
DBLP Embedding ours 0.736$\pm$0.004 0.790$\pm$0.003 0.663$\pm$0.002 0.745$\pm$0.008 0.778$\pm$0.015 0.672$\pm$0.004 0.705$\pm$0.001 0.742$\pm$0.002 0.634$\pm$0.001
Shen et al. 0.835$\pm$0.000 0.911$\pm$0.000 0.826$\pm$0.001 0.817$\pm$0.001 0.910$\pm$0.007 0.818$\pm$0.003 0.832$\pm$0.000 0.913$\pm$0.000 0.827$\pm$0.001
Prediction ours 0.850$\pm$0.000 0.875$\pm$0.001 0.844$\pm$0.001 0.826$\pm$0.007 0.880$\pm$0.007 0.821$\pm$0.005 0.849$\pm$0.001 0.875$\pm$0.001 0.844$\pm$0.009
Shen et al. 0.749$\pm$0.000 0.800$\pm$0.000 0.720$\pm$0.004 0.797$\pm$0.001 0.892$\pm$0.001 0.813$\pm$0.002 0.751$\pm$0.000 0.806$\pm$0.000 0.728$\pm$0.003
Projection ours 0.838$\pm$0.002 0.867$\pm$0.001 0.830$\pm$0.004 0.826$\pm$0.005 0.881$\pm$0.005 0.805$\pm$0.009 0.837$\pm$0.000 0.860$\pm$0.001 0.828$\pm$0.004
Shen et al. 0.844$\pm$0.000 0.911$\pm$0.000 0.837$\pm$0.009 0.827$\pm$0.000 0.911$\pm$0.000 0.822$\pm$0.005 0.838$\pm$0.001 0.931$\pm$0.00 0.833$\pm$0.002
Pubmed Embedding ours 0.851$\pm$0.002 0.867$\pm$0.002 0.851$\pm$0.001 0.841$\pm$0.000 0.874$\pm$0.001 0.831$\pm$0.003 0.845$\pm$0.002 0.847$\pm$0.002 0.851$\pm$0.004
Shen et al. 0.822$\pm$0.001 0.902$\pm$0.000 0.816$\pm$0.005 0.835$\pm$0.001 0.893$\pm$0.001 0.835$\pm$0.001 0.828$\pm$0.000 0.895$\pm$0.001 0.829$\pm$0.004
Prediction ours 0.817$\pm$0.004 0.849$\pm$0.002 0.818$\pm$0.001 0.843$\pm$0.003 0.880$\pm$0.002 0.835$\pm$0.009 0.824$\pm$0.004 0.862$\pm$0.001 0.817$\pm$0.001
Shen et al. 0.682$\pm$0.008 0.713$\pm$0.000 0.671$\pm$0.009 0.742$\pm$0.003 0.787$\pm$0.002 0.721$\pm$0.001 0.624$\pm$0.021 0.661$\pm$0.023 0.655$\pm$0.005
Projection ours 0.730$\pm$0.001 0.749$\pm$0.001 0.821$\pm$0.002 0.757$\pm$0.001 0.783$\pm$0.001 0.736$\pm$0.005 0.695$\pm$0.007 0.712$\pm$0.001 0.663$\pm$0.001
Shen et al. 0.825$\pm$0.001 0.909$\pm$0.000 0.824$\pm$0.009 0.833$\pm$0.001 0.918$\pm$0.002 0.847$\pm$0.002 0.819$\pm$0.001 0.892$\pm$0.001 0.821$\pm$0.005
Citeseer Embedding ours 0.814$\pm$0.006 0.843$\pm$0.005 0.820$\pm$0.002 0.844$\pm$0.001 0.869$\pm$0.006 0.857$\pm$0.009 0.812$\pm$0.005 0.820$\pm$0.005 0.829$\pm$0.002
Shen et al. 0.907$\pm$0.000 0.922$\pm$0.000 0.877$\pm$0.003 0.905$\pm$0.002 0.930$\pm$0.004 0.775$\pm$0.009 0.886$\pm$0.003 0.906$\pm$0.004 0.829$\pm$0.001
Prediction ours 0.922$\pm$0.001 0.932$\pm$0.004 0.905$\pm$0.001 0.896$\pm$0.003 0.925$\pm$0.004 0.902$\pm$0.005 0.879$\pm$0.011 0.895$\pm$0.011 0.868$\pm$0.001
Shen et al. 0.446$\pm$0.002 0.444$\pm$0.000 0.368$\pm$0.009 0.395$\pm$0.091 0.396$\pm$0.086 0.273$\pm$0.002 0.500$\pm$0.002 0.495$\pm$0.002 0.298$\pm$0.009
Projection ours 0.500$\pm$0.006 0.490$\pm$0.006 0.389$\pm$0.009 0.497$\pm$0.008 0.492$\pm$0.008 0.387$\pm$0.001 0.400$\pm$0.075 0.399$\pm$0.071 0.301$\pm$0.003
Shen et al. 0.750$\pm$0.011 0.766$\pm$0.000 0.889$\pm$0.005 0.864$\pm$0.026 0.878$\pm$0.032 0.872$\pm$0.004 0.589$\pm$0.026 0.606$\pm$0.025 0.424$\pm$0.001
Amazon Embedding ours 0.906$\pm$0.005 0.924$\pm$0.003 0.874$\pm$0.004 0.861$\pm$0.032 0.882$\pm$0.034 0.887$\pm$0.004 0.839$\pm$0.011 0.851$\pm$0.012 0.765$\pm$0.009
Shen et al. 0.948$\pm$ 0.001 0.960$\pm$ 0.002 0.943$\pm$0.001 0.906$\pm$ 0.001 0.927$\pm$ 0.002 0.911$\pm$0.003 0.939$\pm$ 0.003 0.953$\pm$ 0.002 0.909$\pm$0.001
Prediction ours 0.948$\pm$0.002 0.959$\pm$0.003 0.944$\pm$0.001 0.895$\pm$0.006 0.911$\pm$0.007 0.918$\pm$0.005 0.944$\pm$0.001 0.956$\pm$0.001 0.921$\pm$0.009
Shen et al. 0.821$\pm$ 0.001 0.827$\pm$ 0.001 0.810$\pm$0.003 0.859$\pm$ 0.002 0.937$\pm$ 0.002 0.850$\pm$0.001 0.799$\pm$ 0.021 0.850$\pm$ 0.031 0.720$\pm$0.003
Projection ours 0.789$\pm$0.006 0.795$\pm$0.007 0.811$\pm$0.005 0.673$\pm$0.020 0.684$\pm$0.021 0.765$\pm$0.002 0.758$\pm$0.009 0.765$\pm$0.008 0.710$\pm$0.001
Shen et al. 0.945$\pm$ 0.001 0.826$\pm$ 0.001 0.944$\pm$0.004 0.920$\pm$ 0.002 0.937$\pm$ 0.001 0.910$\pm$0.004 0.930$\pm$ 0.007 0.947$\pm$ 0.011 0.910$\pm$0.002
Coauthor Embedding ours 0.941$\pm$0.002 0.952$\pm$0.001 0.940$\pm$0.002 0.904$\pm$0.009 0.921$\pm$0.008 0.920$\pm$0.009 0.939$\pm$0.001 0.948$\pm$0.001 0.920$\pm$0.009
Shen et al. 0.898$\pm$0.000 0.947$\pm$0.000 0.891$\pm$0.003 0.900$\pm$0.000 0.958$\pm$0.000 0.886$\pm$0.009 0.892$\pm$0.000 0.936$\pm$0.000 0.889$\pm$0.005
Prediction ours 0.903$\pm$0.004 0.923$\pm$0.005 0.903$\pm$0.009 0.896$\pm$0.007 0.940$\pm$0.003 0.876$\pm$0.001 0.889$\pm$0.012 0.923$\pm$0.007 0.902$\pm$0.002
Shen et al. 0.876$\pm$0.000 0.915$\pm$0.000 0.869$\pm$0.001 0.857$\pm$0.000 0.913$\pm$0.000 0.859$\pm$0.009 0.888$\pm$0.000 0.929$\pm$0.000 0.886$\pm$0.009
Projection ours 0.880$\pm$0.002 0.906$\pm$0.002 0.889$\pm$0.002 0.842$\pm$0.012 0.891$\pm$0.010 0.842$\pm$0.003 0.867$\pm$0.003 0.886$\pm$0.009 0.868$\pm$0.004
Shen et al. 0.900$\pm$0.000 0.927$\pm$0.000 0.897$\pm$0.005 0.845$\pm$0.000 0.880$\pm$0.000 0.892$\pm$0.001 0.839$\pm$0.000 0.855$\pm$0.000 0.840$\pm$0.009
ACM Embedding ours 0.835$\pm$0.010 0.848$\pm$0.016 0.883$\pm$0.001 0.864$\pm$0.013 0.900$\pm$0.011 0.877$\pm$0.003 0.853$\pm$0.003 0.862$\pm$0.007 0.855$\pm$0.003
@@
Accuracy Fidelity F1 score Accuracy Fidelity F1 score Accuracy Fidelity F1 score
Shen et al. 0.761$\pm$0.000 0.904$\pm$0.000 0.682$\pm$0.001 0.759$\pm$0.000 0.864$\pm$0.000 0.689$\pm$0.001 0.764$\pm$0.000 0.891$\pm$0.001 0.683$\pm$0.001
Prediction ours 0.761$\pm$0.002 0.858$\pm$0.002 0.696$\pm$0.001 0.770$\pm$0.007 0.853$\pm$0.005 0.690$\pm$0.001 0.760$\pm$0.005 0.849$\pm$0.008 0.656$\pm$0.001
Shen et al. 0.714$\pm$0.001 0.807$\pm$0.000 0.483$\pm$0.001 0.734$\pm$0.000 0.839$\pm$0.000 0.596$\pm$0.001 0.729$\pm$0.001 0.830$\pm$0.000 0.403$\pm$0.001
Projection ours 0.731$\pm$0.002 0.821$\pm$0.003 0.570$\pm$0.004 0.737$\pm$0.004 0.824$\pm$0.006 0.599$\pm$0.005 0.743$\pm$0.000 0.829$\pm$0.001 0.545$\pm$0.001
Shen et al. 0.757$\pm$0.000 0.887$\pm$0.000 0.637$\pm$0.004 0.776$\pm$0.001 0.885$\pm$0.001 0.710$\pm$0.009 0.762$\pm$0.000 0.895$\pm$0.000 0.626$\pm$0.002
DBLP Embedding ours 0.737$\pm$0.017 0.809$\pm$0.021 0.632$\pm$0.002 0.764$\pm$0.003 0.837$\pm$0.003 0.701$\pm$0.009 0.724$\pm$0.003 0.787$\pm$0.003 0.645$\pm$0.002
Shen et al. 0.855$\pm$0.000 0.956$\pm$0.000 0.851$\pm$0.004 0.839$\pm$0.000 0.944$\pm$0.001 0.840$\pm$0.001 0.856$\pm$0.000 0.952$\pm$0.001 0.847$\pm$0.001
Prediction ours 0.856$\pm$0.002 0.925$\pm$0.001 0.850$\pm$0.003 0.846$\pm$0.003 0.925$\pm$0.003 0.826$\pm$0.009 0.855$\pm$0.001 0.924$\pm$0.000 0.849$\pm$0.004
Shen et al. 0.852$\pm$0.000 0.951$\pm$0.000 0.820$\pm$0.003 0.822$\pm$0.001 0.922$\pm$0.001 0.826$\pm$0.005 0.845$\pm$0.000 0.938$\pm$0.000 0.823$\pm$0.001
Projection ours 0.848$\pm$0.001 0.920$\pm$0.001 0.844$\pm$0.001 0.838$\pm$0.001 0.920$\pm$0.002 0.830$\pm$0.001 0.842$\pm$0.001 0.915$\pm$0.002 0.834$\pm$0.001
Shen et al. 0.859$\pm$0.001 0.958$\pm$0.000 0.853$\pm$0.003 0.842$\pm$0.002 0.944$\pm$0.003 0.838$\pm$0.001 0.848$\pm$0.000 0.953$\pm$0.000 0.847$\pm$0.001
Pubmed Embedding ours 0.859$\pm$0.000 0.920$\pm$0.005 0.854$\pm$0.001 0.848$\pm$0.002 0.924$\pm$0.002 0.844$\pm$0.001 0.847$\pm$0.003 0.895$\pm$0.006 0.847$\pm$0.001
Shen et al. 0.829$\pm$0.001 0.933$\pm$0.000 0.832$\pm$0.005 0.846$\pm$0.001 0.908$\pm$0.000 0.838$\pm$0.001 0.831$\pm$0.000 0.908$\pm$0.000 0.832$\pm$0.009
Prediction ours 0.814$\pm$0.003 0.867$\pm$0.001 0.819$\pm$0.001 0.843$\pm$0.004 0.897$\pm$0.003 0.837$\pm$0.004 0.818$\pm$0.001 0.876$\pm$0.006 0.809$\pm$0.004
Shen et al. 0.685$\pm$0.004 0.727$\pm$0.000 0.617$\pm$0.001 0.736$\pm$0.001 0.792$\pm$0.003 0.734$\pm$0.002 0.719$\pm$0.003 0.756$\pm$0.004 0.602$\pm$0.005
Projection ours 0.718$\pm$0.004 0.762$\pm$0.005 0.707$\pm$0.003 0.750$\pm$0.003 0.797$\pm$0.001 0.734$\pm$0.001 0.666$\pm$0.015 0.704$\pm$0.014 0.655$\pm$0.005
Shen et al. 0.830$\pm$0.001 0.913$\pm$0.000 0.825$\pm$0.004 0.840$\pm$0.001 0.905$\pm$0.001 0.829$\pm$0.003 0.820$\pm$0.002 0.886$\pm$0.001 0.817$\pm$0.001
Citeseer Embedding ours 0.812$\pm$0.005 0.849$\pm$0.004 0.815$\pm$0.001 0.845$\pm$0.002 0.878$\pm$0.001 0.841$\pm$0.001 0.803$\pm$0.004 0.824$\pm$0.002 0.812$\pm$0.004
Shen et al. 0.924$\pm$0.001 0.953$\pm$0.000 0.914$\pm$0.005 0.899$\pm$0.009 0.932$\pm$0.007 0.702$\pm$0.001 0.876$\pm$0.002 0.921$\pm$0.002 0.830$\pm$0.003
Prediction ours 0.922$\pm$0.005 0.932$\pm$0.009 0.916$\pm$0.009 0.909$\pm$0.004 0.930$\pm$0.008 0.857$\pm$0.005 0.899$\pm$0.008 0.913$\pm$0.011 0.867$\pm$0.003
Shen et al. 0.720$\pm$0.004 0.743$\pm$0.000 0.571$\pm$0.004 0.767$\pm$0.011 0.818$\pm$0.008 0.338$\pm$0.002 0.638$\pm$0.002 0.674$\pm$0.002 0.554$\pm$0.001
Projection ours 0.856$\pm$0.013 0.854$\pm$0.014 0.820$\pm$0.009 0.779$\pm$0.010 0.794$\pm$0.004 0.515$\pm$0.002 0.654$\pm$0.028 0.670$\pm$0.033 0.554$\pm$0.001
Shen et al. 0.926$\pm$0.001 0.948$\pm$0.000 0.914$\pm$0.002 0.900$\pm$0.007 0.939$\pm$0.005 0.703$\pm$0.004 0.859$\pm$0.001 0.870$\pm$0.001 0.901$\pm$0.009
Amazon Embedding ours 0.921$\pm$0.003 0.929$\pm$0.003 0.909$\pm$0.005 0.894$\pm$0.018 0.916$\pm$0.035 0.889$\pm$0.005 0.906$\pm$0.004 0.913$\pm$0.005 0.860$\pm$0.001
Shen et al. 0.948$\pm$0.000 0.979$\pm$0.000 0.960$\pm$0.009 0.866$\pm$0.004 0.891$\pm$0.004 0.874$\pm$0.001 0.945$\pm$0.000 0.981$\pm$0.000 0.971$\pm$0.005
Prediction ours 0.946$\pm$0.001 0.967$\pm$0.001 0.961$\pm$0.003 0.884$\pm$0.040 0.909$\pm$0.042 0.898$\pm$0.002 0.947$\pm$0.002 0.971$\pm$0.002 0.973$\pm$0.005
Shen et al. 0.731$\pm$0.001 0.743$\pm$0.000 0.714$\pm$0.005 0.844$\pm$0.007 0.865$\pm$0.008 0.820$\pm$0.001 0.717$\pm$0.005 0.729$\pm$0.005 0.704$\pm$0.004
Projection ours 0.907$\pm$0.011 0.924$\pm$0.011 0.899$\pm$0.004 0.674$\pm$0.053 0.682$\pm$0.054 0.750$\pm$0.001 0.782$\pm$0.003 0.802$\pm$0.005 0.790$\pm$0.001
Shen et al. 0.947$\pm$0.000 0.977$\pm$0.000 0.959$\pm$0.004 0.878$\pm$0.002 0.894$\pm$0.006 0.873$\pm$0.001 0.890$\pm$0.030 0.905$\pm$0.034 0.899$\pm$0.001
Coauthor Embedding ours 0.938$\pm$0.003 0.952$\pm$0.006 0.960$\pm$0.009 0.921$\pm$0.004 0.940$\pm$0.004 0.930$\pm$0.003 0.938$\pm$0.001 0.954$\pm$0.003 0.930$\pm$0.001
Shen et al. 0.907$\pm$0.000 0.950$\pm$0.000 0.896$\pm$0.001 0.902$\pm$0.000 0.950$\pm$0.000 0.888$\pm$0.002 0.905$\pm$0.000 0.943$\pm$0.000 0.893$\pm$0.001
Prediction ours 0.905$\pm$0.003 0.924$\pm$0.002 0.910$\pm$0.009 0.900$\pm$0.006 0.938$\pm$0.002 0.884$\pm$0.001 0.889$\pm$0.006 0.910$\pm$0.012 0.890$\pm$0.001
Shen et al. 0.901$\pm$0.000 0.938$\pm$0.000 0.888$\pm$0.004 0.896$\pm$0.000 0.944$\pm$0.000 0.891$\pm$0.002 0.904$\pm$0.000 0.939$\pm$0.000 0.881$\pm$0.002
Projection ours 0.895$\pm$0.002 0.915$\pm$0.002 0.902$\pm$0.004 0.887$\pm$0.003 0.920$\pm$0.004 0.889$\pm$0.001 0.869$\pm$0.004 0.886$\pm$0.003 0.836$\pm$0.009
Shen et al. 0.907$\pm$0.000 0.952$\pm$0.000 0.905$\pm$0.009 0.868$\pm$0.000 0.919$\pm$0.000 0.881$\pm$0.001 0.870$\pm$0.000 0.895$\pm$0.000 0.867$\pm$0.005
ACM Embedding ours 0.879$\pm$0.004 0.890$\pm$0.007 0.898$\pm$0.009 0.841$\pm$0.027 0.836$\pm$0.033 0.878$\pm$0.009 0.877$\pm$0.007 0.893$\pm$0.006 0.871$\pm$0.009